Of the 6 courses offered by the music department at her college, Kay must choose exactly 2 of them. How many different combinations of 2 courses are possible for Kay if there are no restrictions on which 2 courses she can choose?
Answer by venugopalramana(3286)

Of the 6 courses offered by the music department at her college, Kay must choose exactly 2 of them. How many different combinations of 2 courses are possible for Kay if there are no restrictions on which 2 courses she can choose?THIS IS GIVEN BY
6C2=6!/(2!*(6-2)!=6!/(2!*4!)=6*5/2=15